OpenStudy (anonymous):

for this x2 - 12x + 27 how do you know if its (x+3) or (x-3) the plus and minus confuse me :s

OpenStudy (anonymous):

x1.x2=27 x1+x2=12

OpenStudy (anonymous):

so x1=3 and x2=9

OpenStudy (anonymous):

when you made x=x1or x2 you must get a zero

OpenStudy (anonymous):

so p(x)=(x-x1).(x-x2)

OpenStudy (anonymous):

@RaphaelFilgueiras (x + 3) (x + 9) (x – 3) prime these are the choices it gave me.

OpenStudy (anonymous):

so p(x)=(x-3).(x-9)

OpenStudy (anonymous):

ok thanks c: @RaphaelFilgueiras

OpenStudy (anonymous):

you are welcome